Bryan Carpenter 6 Sandra Close Walsall W. Midlands, WS9 8UW

June 28, 1996

The Lettings Manager, London and Perth, 72 London Road, Southampton, SO15 2AJ

Dear sir,

I vacated the property

Flat 8, Broughton House 28/30 Bullar Road Bitterne Park Southampton, S018 1GS

in January. I enclose proof of payment of final electricity, water and phone bills, and would like to reclaim my surety deposit.

In addition I believe I have overpaid you by one month’s rent. A transfer of £365 was made to your account on 15th January. Can you please also reimburse this sum. Apologies for this oversight.

Yours sincerely,